OpenStudy (anonymous):

@kevinzhengli Solve each equation for y and write the answer in y=mx+b format and the question is 4x=6y-8

OpenStudy (anonymous):

6y=4x+8 y=(4/6)x+(8/6) y=(2/3)x+(4/3)

OpenStudy (anonymous):

What did you do to the 6y? Did you reduce it?

OpenStudy (anonymous):

I divided the left and right by 6 to get rid of it on the left so that the left is only Y.

OpenStudy (anonymous):

So the 4/6 & 8/6 are simplified?

OpenStudy (anonymous):

I'll write it out more clearly.